### v2

# Towards Explainable Fake Image Detection with Multi-Modal Large Language Models

Abstract.
Recent advancements in image generation have provoked social and security concerns, yet most detection methods rely on black-box models that generalize poorly. By utilizing advances in Multi-modal Large Language Models (MLLMs), we propose a framework that fuses six specialized paradigms, each analyzing a distinct aspect of the image, to provide a final verdict with coherent, evidence-based reasoning. Experimental results on a diverse dataset of real and AI-generated images demonstrate that our approach outperforms both traditional detection methods and top humans, while providing explainability. This study underscores the potential of MLLMs in developing robust, explainable, and reasoning-driven detection systems. The code is available at https://github.com/Gennadiyev/mllm-defake.
Generative AI, Multi-modal Large Language Models, Explainable AI, AI safety
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3755421isbn: 979-8-4007-2035-2/2025/10ccs: Computing methodologies Natural language generationccs: Computing methodologies Computer visionccs: Social and professional topics Censorship
1. Introduction
Figure 1. Traditional methods do not provide reasons for their predictions, while our method provides sensible reasons behind the verdict.:
Figure 2. The overall design of the proposed MLLM-based AI-generated image detection framework.:
The rapid emergence of deep learning and artificial intelligence, exemplified by Generative Adversarial Networks (GANs) (Goodfellow et al., 2014) and diffusion models (Dhariwal and Nichol, 2021a), has revolutionized synthetic image generation, creating remarkable opportunities for creative innovation. Yet, these technologies also pose critical societal challenges, particularly in cyber-security, where the proliferation of synthetic images can threaten both individual and public safety (Katirai et al., 2023).
Detecting AI-generated images remains a challenging task, usually formulated as a binary classification problem. While deep learning-based methods have exhibited solid performance (Wang et al., 2020; Tan and Le, 2019), they are prone to overfitting and often rely heavily on large, labeled datasets of real and synthetic images. Alternative approaches focus on statistical anomalies (Ricker et al., 2024; Bi et al., 2023) but frequently struggle with poor generalization, especially when limited to specific generative models or image domains (e.g., facial images). Consequently, there is a pressing need for detection models that not only perform well across diverse generative techniques but also provide interpretability, enabling users to understand why images are flagged.
With the development of multi-modal large language models (MLLMs), researchers have begun exploring their potential in detecting AI-generated images. A key advantage of using MLLMs is their ability to generate predictions with explanations, thus enhancing interpretability. Studies have shown that an open-source MLLM (mPlug-owl2 (Ye et al., 2024)) with 7B parameters can achieve an accuracy of 71.78%, while OpenAI’s GPT-4V reaches 78.03%, outperforming the average human accuracy of 74.51% (Li et al., 2024). Zhang et al. (2024c) highlights the quality of reasons that MLLMs provide when verifying image authenticity. Although these models demonstrate notable potential, MLLMs remain susceptible to hallucination (Gao et al., 2024) and generally cannot surpass the accuracy of traditional methods. Meanwhile, their responses can be sensitive to prompt design (Wu et al., 2024), suggesting the need for more refined techniques to enhance both accuracy and reasoning quality.
In this work, we enhance the capabilities of multi-modal large language models (MLLMs) through an innovative interrogation methodology. By integrating human expertise with MLLM-based reasoning, we design six distinct prompts, each focusing on a distinct aspect of the image. By fusing the model responses from these prompts, we achieve a higher overall accuracy and improved interpretability. Our quantitative and qualitative analysis reveals that MLLMs can process visual information similarly to human perception when distinguishing between real and AI-generated images. The models exhibit strong generalization capabilities, with GPT-4o achieving 93.4% accuracy, outperforming traditional deep learning models (Wang et al., 2020) (91.8%) and AEROBLADE (Ricker et al., 2024) (85.2%), as well as the most accurate human annotator (86.3%).
This framework highlights the promise of MLLMs for robust, transparent, and explainable AI-generated image detection. Moreover, the model can explicitly articulate why an image is classified as generated, a crucial capability for platforms aiming to provide clear explanations to users, as illustrated in Figure 1.
2. Related Works
2.1. AI-Generated Image Detection Methods
AI-generated image detection has received significant attention due to the rapidly improving fidelity of synthetic images produced by Generative Adversarial Networks (GANs) (Goodfellow et al., 2014), autoregressive transformers(Van Den Oord et al., 2017), and diffusion-based models (Rombach et al., 2022; Esser et al., 2024; Song et al., 2020; Ho et al., 2020). Early research primarily focused on artifact detection, extracting discriminative cues from spatial and frequency domains to highlight inconsistencies between real and synthetic images (Zhang et al., 2019; Frank et al., 2020; Chai et al., 2020; Dzanic et al., 2020; Liu et al., 2020). These inconsistencies often stem from upsampling artifacts in GAN pipelines (Frank et al., 2020), texture synthesis mismatches(Liu et al., 2020), or limited high-frequency decay in generated outputs(Dzanic et al., 2020). By learning to detect these subtle traces, models can effectively expose fakes, even when artifacts are visually imperceptible.
Beyond artifact-based methods, deep-learning based methods are also advancing. ResNet and Vision Transformers trained on real and synthetic samples can effectively distinguish real from AI-generated images (Wang et al., 2020; Tan and Le, 2019; Park and Owens, 2024). This paradigm leverages deep networks’ strong feature extraction capabilities to automatically learn discriminative patterns. However, the effectiveness of large-scale training heavily depends on data diversity and quality. Generalization to unseen generative models, particularly from GANs to diffusion-based methods, remains challenging. Meanwhile, as diffusion models produce increasingly photorealistic images, traditional artifact-based cues, including unnatural hues, faulty perspectives, and blurry text, are no longer reliable. Additionally, Grommelt et al. (2024) highlighted that existing models are prone to overfitting on JPEG artifacts, further limiting their robustness.
One branch of complementary research direction focuses on improving model explainability. Most existing detection methods output only a binary classification, failing to indicate how or where synthetic cues are identified. Recent efforts emphasize fine-grained or localized detection, providing more transparent classification rationales. Notable approaches include measuring reconstruction errors between an input image and its diffusion-based reconstruction (Ricker et al., 2024), leveraging multi-branch systems to produce multi-level labels for generated images (Bi et al., 2023), or computing local intrinsic dimensionalities (Lorenz et al., 2023). Some approaches integrate text-image contrastive learning for better explainability (Sundararajan et al., 2017; Simonyan et al., 2013; Selvaraju et al., 2017).
While these methods provide regional clues, they lack human-understandable explanations, limiting their effectiveness in enhancing explainability.
2.2. Multi-Modal Large Language Models
Multi-modal large language models (MLLMs) are large language models (LLMs) enhanced with vision or audio modalities (Lin et al., 2024; Bavishi et al., 2023). Over the past few years, both open-source and proprietary models have significantly advanced vision-language capabilities. Open-source ones include Qwen-VL (Bai et al., 2023; Wang et al., 2024; Qwen Team, 2025) and QVQ (Qwen Team, 2024), InternVL (Chen et al., 2024), Ovis (Lu et al., 2024b), LLaVA and its predecessors (Liu et al., 2023b, a, 2024), Llama-3 (The Llama Team, 2024), among many others. Proprietary models have also progressed rapidly. Notable releases include GPT-4o, GPT-4o-mini (OpenAI, 2024b, a) and Gemini-1.5 (Gemini Team, 2024b, a). These models demonstrate remarkable performance in vision-related tasks, including image reasoning and interpretation. Most benchmarks for vision-language models focus on reasoning capabilities related to the image (Ying et al., 2024; Liao et al., 2024) or geometric problem-solving skills (Kamoi et al., 2024; Lu et al., 2024a).
Li et al. (2024) proposed FakeBench, a benchmark for general AI-generated image detection that includes three primary datasets: reasoning, interpretation, and open-question answering. It covers ten generative models across diverse image content, revealing that current MLLMs demonstrate moderate detection accuracy, basic reasoning abilities, and limited fine-grained authenticity analysis. Zhang et al. (2024c) created A-Bench, which “challenges both high-level semantic understanding and low-level visual quality perception” of MLLMs in the form of multiple-choice questions, pointing out major flaws in each AI-generated image. The question-answer pairs are curated by humans and not actively generated by MLLMs. FakeReasoning (Gao et al., 2025) conducts the forgery reasoning task in structured steps using MLLMs, giving promising results by fine-tuning LLaVA-1.5-13b on the MMFR-Dataset, created using GPT-4o and human experts.
Our proposed method significantly enhances the accuracy while also providing fine-grained explanations of why an image is considered real or generated, addressing these limitations.
3. Methodologies
Given an image $x$, we employ a detection function $f$ to classify it as either real (i.e., natural) or fake (i.e., AI-generated), producing a prediction $y=f(x)$. Our approach is driven by the observation that different prompt formulations enable MLLMs to focus on distinct aspects of an image.
Techniques such as image cropping (Zhang et al., 2024b), few-shot learning (Zhao et al., 2023; Guo et al., 2024), and region-of-interest selection (Wu and Xie, 2023) enhance MLLM performance by refining visual feature extraction and interpretation. These techniques influence how models extract and analyze relevant image attributes, leading to more effective detection. Prompt formulation plays a critical role in MLLM performance. We design optimized prompts that effectively integrate textual and visual information, mimicking human perception. By combining domain expertise with MLLM reasoning, our approach enhances both explainability and accuracy. The proposed framework is illustrated in Figure 2, and the following sections provide a detailed explanation of our prompt construction and detection pipeline.
3.1. Prompt Engineering
We provide visual images and textual prompts to the MLLMs, leveraging chain-of-thought (CoT) prompting to improve reasoning. Our approach structures MLLM sessions (i.e., context window) as sequential queries, where each query is defined as: $o_{i}=M(t_{i},v_{i},c_{i-1})$, where $i$ denotes the current query, $M$ represents the MLLM, $o$ is the model’s output, $t$ and $v$ are text/visual prompts (image type in this task) from the user, and $c$ is the session context (query history) from previous queries. Each query triggers a computational payload, incurring processing costs. The output $o$ is incorporated into the next query as part of the updated context: $c_{i+1}=(t_{i},v_{i},o_{i},c_{i})$. To ensure consistency and mitigate randomness, we utilize a predefined $o$ as an assistant reference, summarized by the user. This approach also optimizes inference efficiency by reusing previous outputs, improving model stability and reducing computational costs while maintaining fast inference speed.
Verdict Prompt (P0).
In our method, a designated verdict prompt $n$ is introduced, where the input image $v_{n}$ corresponds to $x$, and the prompt $t_{n}$ is structured as follows:
[Is this image real or fake? End your response with either "real" or "fake".]
This enforces a concise model prediction, allowing us to directly use the output $o_{n}$ as the final classification result $y$. P0 serves as the baseline of our framework, leveraging MLLMs’ classification capabilities. Unless otherwise specified, our approach always concludes with P0.
To systematically explore prompt variations, we develop six distinct prompt strategies, categorized into three types: General Prompts (direct analysis mimicking human reasoning), Few-Shot Prompts (providing a few examples to guide model predictions) and Content-Based Prompts (first analyzing image content before predicting). We briefly introduce each prompt type later. For full implementation details, please refer to the provided code.
System Prompt.
To enhance the interoperability of our method, we define an initial system prompt $h_{0}$ that defines the session’s general settings:
[You are an AIGC detection specialist. The user will ask you about whether an image is real or generated. Observe the image carefully to decide whether it is real or generated. Explain your reasons, and end your response with either "real" or "generated".]
This ensures that the model justifies its classification decision, maintaining explainability in the prediction function $f$. For all subsequent prompts P1–P6, this system prompt initializes the session.
3.2. General Prompts
Humans analyze images top-down, first grasping the overall scene before focusing on key objects and details. We design our model to follow a similar process, leveraging MLLMs’ reasoning abilities to detect anomalies. Our framework systematically analyzes primary objects, assessing their properties against real-world constraints and logical coherence. We design three prompts: Defect Query (P1): Conducts a broad inspection to identify general anomalies; (P2): Focuses on specific regions to refine local anomaly detection; and Common Sense Reasoning (P3): Applies logical reasoning to verify plausibility against real-world knowledge. By integrating these perspectives, our approach enhances the reliability and explainability of AI-generated image detection.
Defect Query (P1).
Generative models rely on extensive real-world data, yet no dataset can fully encapsulate all causal factors of a scene. Photography follows a cause-and-effect paradigm—objects are shaped by lighting, camera angles, lens parameters, and the photographer’s intent. Models trained solely on photographic data often fail to internalize these relationships, leading to artifacts such as incorrect shadows, unnatural object placements, and abnormal focus or blur. To address this, the defect query prompt directs the MLLM to identify such anomalies. The model is informed that the image may contain defects and is tasked with listing potential issues. The process consists of three queries. In the first two queries, $o_{1}$ lists known AI-generated artifacts, $o_{2}$ highlights features typical of real images, and $t_{1}$ / $t_{2}$ are
[What are common defects in AI-generated images?] / [What are the features that real images often have?]
Finally the model makes a final prediction with P0, reasoning based on $o_{1}$ and $o_{2}$. This query structure enables MLLMs to predict and justify their classification by identifying defects such as abnormal object proportions, unrealistic reflections, etc.
Regional Analysis (P2).
Each input image contributes a fixed number of tokens to the MLLM. To emphasize key areas, this method extracts Regions of Interest (ROI) from the original image, allowing the model to focus on specific objects or sub-scenes, such as facial features or text on a book. While this approach may omit broader context, it enhances detail analysis in a controlled setting—especially useful when the main object occupies only a small portion of the frame. This method is inspired by prior work on synthetic image detection Simonyan et al. (2013), where heatmaps are used to improve model explainability. The process consists of two queries. In the first query, $v_{1}=x$, $o_{1}$ will request user confirmation of the ROI, and $t_{1}$ is
[Instead of analyzing the full image, focus on the ROI in the next image to determine whether it is real or generated.]
We extract the top three most relevant regions using DINOv2 (Oquab et al., 2024), generating heatmaps that are thresholded to form bounding boxes from binary pixel maps. Pretrained weights from ImageNet are used. For the second query, $v_{2}$ is the ROI image and $t_{2}$ is
[Here is the ROI. Please indicate whether the image is real or generated based on ROI. End your response with either "real" or "generated"].
The final prediction is then generated, along with an explanation focusing on the main area of the image. This ensures that the model prioritizes fine details of key objects, such as pattern distortions, unrealistic textures, and other detail problems, enhancing detection accuracy.
Common Sense Reasoning (P3).
AI-generated images often contain logical inconsistencies that are implausible in the real world, such as incorrect numbers of fingers, unnatural object attachments, and distorted text. While humans easily recognize these anomalies, MLLMs may overlook them when prioritizing overall image realism. To address this, and inspired by (Zhang et al., 2024a), we prompt the MLLM to detect inconsistencies by comparing images against real-world patterns. This method uses a single query with $v_{1}=x$ and $t_{1}$ is
[ Please analyze this image for any violations of common sense or real-world logic. Consider: 1. Physical proportions (e.g., number of fingers, limbs); 2. Spatial relationships between objects, such as merged textures; 3. Natural world physics and rules, such as unnaturally bent fences; 4. Other logical inconsistencies in the scene Explain any inconsistencies you find. If you find many unexplainable inconsistencies, the image is likely generated.]
This approach guides the model to examine critical visual cues, such as limb counts and spatial relationships. While these anomalies may be subtle, they serve as strong indicators of AI-generated content. By incorporating logical consistency checks, this method improves model generalization across diverse images while enhancing interpretability in “generated” classification.
3.3. Few-Shot Prompt
Few-Shot (P4).
In the realm of MLLM, few-shot learning refers to the technique of providing exemplars in the prompt before the query. Our few-shot prompt provides two labeled exemplars along with their human-annotated responses to a baseline prompt, refining the model’s conceptual understanding. Our prompt design consists of three turns, each using a structured text prompt similar to P0 for prediction. The first two queries introduce labeled exemplars, one real for $v_{1}$ and one fake for $v_{2}$, with manually annotated justifications for their classification. These responses replace $o_{1}$ and $o_{2}$, and are packaged with previous images in a single history payload. Finally, P0 will be conducted for the final prediction. This method leverages the MLLM’s reasoning and generalization capabilities under data constraints, ensuring interpretability. By referencing only a few exemplars, the model’s decision process remains transparent and grounded in direct comparisons. The compact dataset encourages the model to focus on essential features, such as object outlines, textures, and logical color consistency, instead of relying on memorized distributions, leading to more robust conclusions.
3.4. Content-Based Prompts
Content-Based Prompts evaluate an image’s internal coherence by analyzing its main subject, structural patterns, and alignment with real-world expectations. These prompts leverage the MLLM’s ability to identify key features and provide interpretable justifications for inconsistencies. To prevent conflicts between different queries, the process follows a structured multi-session approach. The first session is Object Identification, where $t$ is
[What is the main subject in the image? Please answer concisely with a short descriptive noun phrase, e.g., ’a red bus’, ’a daisy’.]
The second session then builds upon the response from the first turn, refining the analysis with $v=x$. $o$ will be the {class} of $x$. Then the {class} information is integrated into the following sessions for the final prediction, which will be detailed in the next paragraphs.
Structural Analysis (P5).
This method systematically identifies internal inconsistencies within an image based on its object class. The detection process follows a two-step query approach: 1. Component Identification, where $t$ is
[This image shows a {class}. List its key components before determining whether it is real or generated.]
$v=x$, and $o$ outputs a list of expected components based on the identified object class. 2. Structural Verification: The model checks for missing or misplaced components, assessing whether these structural faults indicate AI generation. At the same time, P0 is applied to finalize the classification. This modular approach enhances interpretability by mapping each detected anomaly to a logical explanation, conditioned on the expected class structure. Using this prompt, the model can effectively detect AI-generated images with: unnaturally uniform repetitions (e.g., identical windows in a building, consistent distortions across a fence); absence of critical features (e.g., a dog without a tail, a person missing hands); implausible positioning of objects (e.g., a tree growing out of a building, a car parked on a rooftop).
Stereotype Matching (P6).
This method first classifies the dominant object in an image, then evaluates whether its features conform too closely or uniformly to common stereotypes. Examples include excessively symmetrical human faces, unnaturally uniform animal textures, or exaggerated stereotypical attributes (e.g., a red-beaked rooster with atypical pigmentation). The detection will first conduct stereotype analyses session, where $t$ is
[If an image shows {class}, what are some common stereotypes or patterns you would expect? List at least three.]
$o$ outputs common stereotypes of the identified class. Then, we start a new session. The model analyzes the image for the presence of these common stereotypes, explaining any anomalies and determining whether the image is real or AI-generated with P0. A certain degree of stereotypical accuracy enhances realism, but excessive adherence to these patterns may indicate AI generation, reflecting biases learned from training data. By prompting the MLLM to assess how closely an image follows stereotypical features, this method enhances interpretability in detecting unnatural uniformity or exaggerated regularities that deviate from real-world variation.
3.5. Fusion Process
Prompts categorized as General, Few-Shot, and Content-Based enhance AI-generated image detection via MLLMs by targeting interpretability aspects like local inspections (P2, P4), logical consistency (P3, P5), and real-world traits (P1, P6). We propose two fusion methods to determine the final verdict. The primary approach executes P1–P6 in parallel before consolidating an output, balancing accuracy and explainability. Alternatively, a majority vote method tallies individual prompt verdicts, trading some explainability for computational efficiency while remaining effective in most cases. Experimental results validate these trade-offs, showing the sequential method excels in interpretability and the majority vote offers practical efficiency.
4. Experiments
4.1. Experimental Setup
Prompt GPT-4o GPT-4o-mini Llama 3.2 VI LLaVA-CoT
All Real Generated All Real Generated All Real Generated All Real Generated
P0 0.852 0.855 0.849 0.769 0.910 0.628 0.561 0.645 0.477 0.594 0.768 0.420
P1 0.846 0.802 0.888 0.776 0.972 0.579 0.561 0.645 0.477 0.592 0.749 0.435
P2 0.848 0.824 0.871 0.776 0.846 0.706 0.595 0.637 0.552 - - -
P3 0.896 0.922 0.869 0.701 0.721 0.680 0.531 0.412 0.650 0.600 0.619 0.580
P4 0.834 0.801 0.867 0.860 0.874 0.846 0.497 0.525 0.468 0.526 0.583 0.469
P5 0.841 0.813 0.868 0.734 0.732 0.735 0.497 0.297 0.697 0.610 0.752 0.467
P6 0.888 0.880 0.895 0.753 0.767 0.738 0.584 0.563 0.604 0.586 0.603 0.568
w/o P1 0.895 0.917 0.872 0.829 0.862 0.796 0.605 0.512 0.698 0.669 0.685 0.653
w/o P2 0.863 0.850 0.876 0.790 0.823 0.757 0.609 0.534 0.683 0.674 0.719 0.628
w/o P3 0.871 0.850 0.891 0.799 0.835 0.763 0.628 0.603 0.653 0.667 0.706 0.627
w/o P4 0.913 0.923 0.903 0.770 0.851 0.689 0.612 0.523 0.700 0.690 0.725 0.654
w/o P5 0.865 0.848 0.882 0.786 0.856 0.715 0.623 0.580 0.666 0.665 0.683 0.647
w/o P6 0.875 0.874 0.875 0.765 0.826 0.704 0.587 0.515 0.658 0.677 0.720 0.633
Maj. 0.925 0.949 0.901 0.878 0.895 0.861 0.634 0.653 0.615 0.697 0.769 0.625
Fusion 0.934 0.955 0.912 0.876 0.893 0.859 0.639 0.660 0.617 0.696 0.767 0.625
AEROBLADE CNNSpot CommunityForensics ObjectFormer
All Real Generated All Real Generated All Real Generated All Real Generated
- 0.852 0.834 0.870 0.918 0.978 0.857 0.861 0.850 0.872 0.903 0.910 0.815
DMimageDetection NPR Human (Average) Human (Best)
Public Real Generated All Real Generated All Real Generated All Real Generated
- 0.871 0.924 0.870 0.787 0.915 0.660 0.819 0.848 0.790 0.863 0.910 0.815
Table 1. Quantitative Comparison Results on Accuracy. The fusion of P1-6 is giving the best result across all MLLMs, with GPT-4o surpassing traditional methods and the best human on the task. The Maj. row shows the majority vote result.
We ensure relevance and effectiveness in real-world AI-generated image detection by covering a diverse range of generation models, including but not limited to GANs, auto-regressive models, and diffusion models. To achieve this, we selected 1,000 real images and 1,000 AI-generated images from WildFake (Hong et al., 2025), spanning different generation architectures, with detailed distribution provided in the appendix. To benchmark MLLM performance against humans, we invited 24 volunteers to classify the images, 12 of whom had prior experience with AI image generation.
We tested recent MLLMs such as GPT-4o, GPT-4o-mini, Llama 3.2 VI (Llama-3.2-Vision-Instruct 11B) (The Llama Team, 2024), LLaVA-CoT (Xu et al., 2024), Qwen-VL (Bai et al., 2023; Wang et al., 2024; Qwen Team, 2025, 2024), Ovis (Lu et al., 2024b) and InternVL series (Chen et al., 2024). Among them, we selected four models for our full analyses: GPT-4o, GPT-4o-mini, Llama 3.2 VI, and LLaVA-CoT. In our experiments, we used OpenAI GPT-4o that points to gpt-4o- 2024-08-06 and GPT-4o-mini that points to gpt-4o-mini- 2024-07-18. The Llama 3.2 VI refers to the Llama-3.2-Vision-Instruct 11B model, from which LLaVA-CoT is a fine-tuned chain-of-thought model. Local MLLMs are deployed using vllm-0.7.2 (Kwon et al., 2023) on four NVIDIA A100-40G GPUs.
To compare the performance with traditional methods, we select AEROBLADE (Ricker et al., 2024), CNNSpot (Wang et al., 2020), CommunityForensics (Park and Owens, 2024) and ObjectFormer (Wang et al., 2022). We evaluate them on the same sampled dataset used for MLLM and human evaluations. For AEROBLADE, we employ Stable Diffusion versions 1.5, 2.1, and 3.5-large (Rombach et al., 2022; Esser et al., 2024) as reconstruction models and select SD3.5-Large for the results, since it provides the best performance. For CNNSpot and ObjectFormer, since the pretrained models performed suboptimally, we evaluated on checkpoints that we trained on the same data. CommunityForensics requires no training.
Figure 3. (a) Examples where GPT-4o fusion gives correct result, while CNNSpot, AEROBLADE, and vanilla GPT-4o (P0) fail. (b) A fusion example. GPT-4o can combine responses from P1-6 effectively, drawing conclusions from reasons instead of verdicts.:
Time (s.) GPT-4o GPT-4o-mini Llama 3.2 VI LLaVA-CoT
P0 1.4 0.9 3.2 3.4
P1-6 47.9 25.0 67.7 69.1
P1-6 + F. 52.2 29.6 78.1 81.5
P1-6 (Parellel) 11.4 7.2 32.8 34.0
P1-6 (Parellel) + F. 16.2 6.8 37.9 41.3
Table 2. Evaluation time for different models and different prompts, in seconds per image. The time reported includes image processing.
Table 2 reports the average inference time per image when six prompts are queried both sequentially and with parallelism enabled. For Llama 3.2 VI (The Llama Team, 2024) and LLaVA-CoT (Xu et al., 2024), profiling was conducted on 8× NVIDIA A100-40G GPUs. Network latency is included for models relying on OpenAI API calls (OpenAI, 2024b, a). Results indicate that P5 and P6 require more time, as they involve larger, multi-turn payloads, but all models maintain reasonable inference times with parallelism, given that natural language output remains rate-limited by token throughput of MLLMs.
4.2. AI-Generated Image Detection Accuracy
The results are summarized in Table 1, reporting accuracy for MLLMs: GPT-4o, GPT-4o-mini, Llama 3.2 VI, LLaVa-CoT; traditional methods: AEROBLADE (Ricker et al., 2024), CNNSpot (Wang et al., 2020), CommunityForensics (Park and Owens, 2024), DMimageDetection (Corvi et al., 2023) and NPR (Ojha et al., 2023); and human performance. For MLLMs, we report results for baseline (P0), all proposed prompts, majority vote of P1-P6 (Maj.), and fused results. Humans performed well overall (81.9% accuracy), particularly on real images (84.8% accuracy). The best volunteer achieved 4.4% higher accuracy than the average human performance and slightly outperformed the best MLLM baseline (86.3% vs. 85.2% for GPT-4o P0). Other MLLMs performed worse than GPT-4o. While CNNSpot (91.8%) surpassed GPT-4o (85.2%), it operates as a black-box model.
Not all P1–P6 outperform P0, which is expected since these prompts prioritize both accuracy and interoperability. This is a trade-off rather than a direct improvement in both aspects. As a scaled-down version of GPT-4o, GPT-4o-mini performs significantly worse across prompts. They have different trends with different prompts. GPT-4o excels in P3 (reasoning) and P6 (stereotype matching). GPT-4o-mini struggles the most on P3, while in P1 (defect detection), GPT-4o outperforms GPT-4o-mini, particularly in detecting generated images.
A key finding is the performance gain from response fusion. When using majority vote from verdicts provided by P1 P6, we consistently achieve better accuracy over independent prompts. To fully utilize previous responses, when fusing together for a final verdict, GPT-4o and Llama 3.2 Vision Instruct show an increase in accuracy. While the other two models did not benefit from this process, the aggregated reasons provided by the fusion process are more sensible than individual outputs. When fused, GPT-4o can reach 93.4% accuracy, surpassing CNNSpot. Both majority voting and fusion give more accurate results when compared to P0: Fused compared to P0, GPT-4o +8.2, GPT-4o-mini + 10.7, Llama 3.2 VI + 7.8, LLaVA-CoT + 10.4.
These results align with prior research, demonstrating that model ensembling enhances generalization, accuracy, and reduces false positives and false negatives in MLLM-based AI-generated image detection.
Prompt GPT-4o GPT-4o-mini Llama 3.2 VI LLaVA-CoT
Accuracy Rejections ($\downarrow$) Accuracy Rejections Accuracy Rejections Accuracy Rejections
P0 (‘Generated’) 0.852 26 0.769 23 0.561 6 0.592 5
P0 (‘Fake’) 0.849 45 0.761 49 0.543 8 0.588 5
Table 3. Changing ‘fake’ to ‘generated’ in the prompt can lower rejection rates and improve accuracy for all models.
4.3. Prompt Effectiveness
Figure 4. Percentage of cases where prompts give different verdicts. The data is aggregated from all four models evaluated.:
To assess the effectiveness of our six prompts, we calculated the probability of divergent verdicts across the prompts for a given image, with results presented in Figure 4. Among all real and generated samples, 22.31% received at least one contradictory verdict across prompts from GPT-4o, and 31.44% from LLaVA-CoT. This demonstrates that different prompts can guide the model to focus on different aspects of the image, leading to independent, different decisions.
We also conducted an ablation study by systematically excluding each prompt from the fusion process while preserving the reasoning structure. By analyzing the resulting accuracy, we assess the significance of each prompt in enhancing the model’s decision-making process. If a prompt is redundant or ineffective, its removal should have minimal impact on the fused accuracy. Conversely, if a prompt is critical, omitting it should lead to a noticeable accuracy drop. For each iteration, we remove one prompt at a time and re-run the fusion queries, comparing the results against the original fused accuracy reported in  Table 1. Different models exhibited different sensitivities to prompt removal. For example, GPT-4o and GPT-4o-mini showed the least accuracy drop when omitting P1 (defect detection), whereas open-source models retained more accuracy when ablating P1. No ablation configuration outperformed the full fusion model (P1–P6), demonstrating the effectiveness of all proposed prompts.
4.4. Qualitative Results on Reasoning
Figure 3 (a) presents two cases where traditional methods and P0 fail, but fused GPT-4o correctly classifies the images by integrating regional observations and general cues. In the first case, GPT-4o with fused prompts identifies realistic skin texture, consistent lighting, genuine facial expression, and naturally flowing hair, leading to a real classification. In the second case, GPT-4o with fused prompts detects overly smooth facial details, unnatural shadows, disproportionate background elements, and poor integration, resulting in a fake classification. These cases demonstrate GPT-4o’s holistic reasoning, effectively combining multiple perspectives for accurate detection.
A key advantage of our pipeline is the MLLM fusion stage, which corrects errors from individual prompts (P1–P6). As shown in Figure 3 (b), while P5 incorrectly classifies an image as real, the model re-evaluates observations from P3 and P6, concluding that the image is AI-generated. By integrating previous responses, the model enhances explainability, reinforcing final classifications with well-structured reasoning.
4.5. Discussions on MLLM Rejections
MLLMs may generate non-parsable responses or refuse to comment on the given image, termed “rejections,” when processing certain images. The rejection rate varies across different models, and the types of images rejected differ accordingly. A frequent justification for these rejections by MLLMs is their inability to provide commentary on images, particularly those depicting faces, due to ethical considerations. These rejections can be mitigated by modifying the prompt, leading to lower rejection rates. Table 3 shows that among the 2,000 images evaluated, replacing the word “fake” with “generated” resulted in a 0.95% (19 samples) reduction in rejections for GPT-4o and 1.3% for GPT-4o-mini. In comparison, open-source models generally exhibit lower rejection rates. Open-source models such as Llama-3.2-Vision-Instruct and LLaVA-CoT have a much lower rejection rate than GPT-4o and -4o-mini.
5. Limitations
AI-generated content introduces significant ethical risks, particularly in misinformation, identity manipulation, and fraud. Robust detection systems are essential to address these concerns. Ethically, ensuring transparency and accountability in detection models is critical, especially in sensitive areas like forensics and law enforcement. These models must be interpretable and free from biases that could lead to unfair outcomes. Privacy and security concerns also arise from the use of AI in surveillance and public discourse, requiring strict safeguards.
While MLLMs show promise in detecting AI-generated images, challenges remain in interpretability and alignment with human perception. Further research is needed to quantify this alignment and assess how human evaluators interpret MLLM-generated explanations in forensic applications. Future research could explore adaptive fusion mechanisms that dynamically adjust reasoning strategies based on image characteristics, as well as adversarial training on high-quality AI-generated images to enhance model robustness.
6. Conclusion
In this work, we explore the use of MLLMs for AI-generated image detection, offering a human-interpretable classification method. We introduce six detection paradigms tailored for MLLMs and validate their effectiveness across diverse image generator architectures. A majority vote from these paradigms can surpass traditional classification methods and the most accurate humans. The fusion of model responses, when passed again to the MLLM, allows for aggregated and accurate reasoning of why an image is considered real or AI-generated. Combining diverse heuristics and reasoning improves both accuracy and reliability. Our ablation study confirms that leveraging multiple detection paradigms outperforms single-method classification.
Acknowledgements.
This research is supported, in part, by the National Natural Science Foundation of China (Grant No. 62302295), by Ant Group, the Shanghai Municipal Science and Technology Major Project, China (Grant No. 2021SHZDZX0102), the Pioneer R&D Program of Zhejiang Province (No. 2024C01024) and the foundation of Key Laboratory of Artificial Intelligence, Ministry of Education, P.R. China.
Appendix
Prompt GPT-4o GPT-4o-mini Gemini 1.5 Pro Kimi-VL-A3B-Instruct
All Real Generated All Real Generated All Real Generated All Real Generated
P0 0.852 0.855 0.849 0.769 0.910 0.628 0.758 0.906 0.610 0.585 0.622 0.548
MiMo-VL-7B Qwen2-VL-7B-Instruct QVQ-72B-Preview Qwen2.5-VL-7B-Instruct
All Real Generated All Real Generated All Real Generated All Real Generated
P0 0.585 0.887 0.283 0.623 0.901 0.346 0.643 0.789 0.497 0.621 0.660 0.582
Ovis1.6-Gemma2-9B InternVL2-8B-MPO InternVL2.5-26B MiniCPM-V-2.6 (8B)
All Real Generated All Real Generated All Real Generated All Real Generated
P0 0.559 0.828 0.290 0.592 0.623 0.561 0.566 0.673 0.460 0.545 0.766 0.324
Table 4. Experimental result of P0 on different open-source and proprietary models.
Appendix A Dataset Distribution
Figure 5. Distribution of AI generation methods among images in the dataset.:
To ensure diversity, we curated a dataset containing 2000 images from multiple sources, covering three major architectures and variations of specific models, as detailed below.
Diffusion-Based Generators. We sampled from Midjourney (Holub, 2022), SDv2 (Rombach et al., 2022), ADM (Dhariwal and Nichol, 2021b), DDIM (Song et al., 2020), DDPM (Ho et al., 2020), Imagen (Saharia et al., 2022), VQDM (Gu et al., 2022), ControlNet (Zhang et al., 2023), SD with adaptors (LoRA (Hu et al., 2021), LyCORIS (Yeh et al., 2023)) and DALLE 3 (OpenAI, 2023).
GAN-Based Generators. We sampled from DF-GAN (Tao et al., 2022), GALIP (Tao et al., 2023), StyleGAN (Karras et al., 2018), VQGAN (Esser et al., 2021), GigaGAN (Kang et al., 2023), BigGAN (Brock et al., 2018) and StarGAN (Choi et al., 2018, 2020).
Other Generators. We sampled from VQVAE (Van Den Oord et al., 2017), MAGE (Li et al., 2023) and MAE (He et al., 2022).
Real Images. We select real images from COCO (Chen et al., 2015), LAION-5B (Schuhmann et al., 2022), LSUN Church (Yu et al., 2015), FFHQ (Karras et al., 2018), CelebA-HQ (Karras et al., 2017), ImageNet (Deng et al., 2009) and AFHQ (Choi et al., 2020).
Appendix B Discussions on Model Responses
To accommodate paper length constraints, we present additional results for qualitative analysis in Figure 9. These examples are unfiltered, showcasing the model’s typical performance. Our analysis shows that GPT-4o produces varying verdicts across different prompts but maintains consistent reasoning accuracy during the fusion stage. This aligns with our experimental findings, where post-fusion responses achieve the highest accuracy among MLLM-based approaches. Our full prompts can be accessed from this link.
For conciseness, responses are further shortened with the prompt below.
<original_response>
Summarize the above response into no more than 5 key points, where each point should be fewer than 6 words. Output a numbered list in markdown format. Do not include the final verdict in your response.
Figure 6. Common key words present in model responses.:
Key words that model uses. To qualitatively assess the model’s attention tendency, we list the most common key words that are seen many times in model responses. While some of the model responses
Appendix C Experimental Results on Other MLLMs
We evaluate several MLLMs to distinguish AI-generated images from real ones. The results, as shown in Table 4, indicate that Qwen2-VL (Wang et al., 2024) and QVQ (Qwen Team, 2024) perform relatively well in this task. However, Qwen2-VL exhibits a bias towards classifying images as real, while QVQ, with 72B parameters, can take almost ten minutes for one fusion query on 4x NVIDIA H800 GPU. Qwen2.5-VL-7B-Instruct also gives remarkable results, but it lacks the multi-turn image input chat template, rendering P2 and P4 unusable. Gemini 1.5 Pro (Gemini Team, 2024a), as a proprietary model, scored lower than GPT-4o-mini in all three metrics. We selected LLaVA-CoT (Xu et al., 2024) for its reinforced CoT capabilities and conducted further experiments to explore the potential of open-source MLLMs in AI-generated image detection. Llama 3.2 Vision Instruct, the base model of LLaVA-CoT, is used for comparison.
Figure 7. Grid representation of accuracy changes for different real-fake exemplar pairs in the few-shot learning setting. Positive values indicate accuracy improvements over the zero-shot baseline (P-Zero), while negative values denote declines.:
Figure 8. Impact of real and fake sample selection on the overall accuracy of GPT-4o. The dashed horizontal line score represents zero-shot accuracy, while the solid horizontal line line denotes the average accuracy across all two-shot pairs.:
Appendix D Discussions on Few-Shot Exemplar Selection
MLLMs exhibit sensitivity to exemplar selection in few-shot learning (Guo et al., 2024). To systematically assess its impact, we compare few-shot learning (P3) against the zero-shot paradigm (P0), evaluating how exemplar choice affects model accuracy and decision stability.
We selected six real and six AI-generated images uniformly. These images were independently annotated by human researchers, identifying artifacts, inconsistencies, and perceptual anomalies that might influence MLLM classification. All 36 possible real-fake combinations were tested on the same dataset as used in Section 4. Additionally, we removed both exemplars from the prompt to create a zero-shot context (P-Zero), serving as the control group.
Figure 8 presents the comparative accuracy of different few-shot sample combinations against the zero-shot baseline. Additionally, Figure 7 provides a grid representation of accuracy variations across different pairings of real and fake exemplars.
From the results, we observe that model accuracy fluctuates depending on the selection of exemplars. The highest observed accuracy gain was +1.60%, while the largest drop was -1.00%. Despite these variations, the overall standard deviation remains moderate, indicating that while sample selection plays a role, GPT-4o exhibits relative stability in performance across different trials.
While sample selection influences performance, the variance remains limited, suggesting GPT-4o exhibits robustness to exemplar variations. In applications demanding high interpretability and reliability, exploring broader sample combinations could further optimize performance.
Figure 9. More examples with GPT-4o. Responses are summarized. Different colors of the text indicate the verdict for each query (red/green: fake/real).: